# A Framework for Holistic KLD-based Waveform Design for Multi-User-Multi-Target ISAC Systems

### v4

@@ X Conclusion
This paper has proposed and validated a novel KLD-based framework for optimising ISAC systems. The three considered optimisation techniques, namely, K-ROP, K-COP, and K-IOP for ISAC approach, each demonstrated significant improvements over the non-optimised scenario. K-ROP yielded substantial enhancements in both target detection and communication performance while maintaining efficient computational requirements, as evidenced by the numerical results. K-COP, while primarily benefiting the communication subsystem, also showed modest improvements in radar performance, particularly at lower $P_{\mathrm{R}}/N_{o}$ levels. Both techniques exhibited robust performance across varying SNR values. In addition, the K-IOP effectively unifies these advantages, consistently outperforming the two separate approaches in terms of detection probability, BER, and estimation accuracy. By jointly designing radar and communication waveforms, the K-IOP approach leverages the synergy between sensing and data transmission, thereby achieving higher overall KLD and a more balanced performance for future ISAC networks. Moreover, the numerical results revealed that exploiting the communication signal as part of the radar subsystem significantly boosts ISAC system capabilities, underscoring the importance of holistic design in shared-antenna deployments. These findings provide valuable insights for designing and optimising ISAC systems, offering a promising pathway to balance sensing and communication requirements in next-generation wireless networks. The proposed framework offers the flexibility to adapt system resources for diverse operating regimes, paving the way for more efficient and versatile wireless communications and sensing technologies.
Appendix A Gradient Derivations
This appendix provides detailed derivations for the gradients used in our optimisation problem. The objective function $f(\mathbf{\dot{W}}_{\mathrm{r}})$ is given in (31), where the $\mathrm{KLD}_{\mathrm{r},t}$ is defined in (27). Therefore, to derive $\nabla f(\mathbf{\dot{W}}_{\mathrm{r}})$, we first compute $\partial\mathrm{KLD}_{\mathrm{r},t}/\partial\mathbf{W}_{\mathrm{r},t}$. Using the matrix identity $\frac{\partial\ln(\det(\mathbf{X}))}{\partial\mathbf{X}}=(\mathbf{X}^{-1})^{T}$, and derivative $\frac{\partial\mathrm{Tr}(\mathbf{X}^{-1}\mathbf{A})}{\partial\mathbf{X}}=-(\mathbf{X}^{-1})^{T}\mathbf{A}^{T}(\mathbf{X}^{-1})^{T}$ we obtain,
$\frac{\partial\ln(\det(\mathbf{R}_{2,t}))}{\partial\mathbf{R}_{2,t}}\!=\!\mathbf{R}_{2,t}^{-1},\quad\frac{\partial\mathrm{Tr}(\mathbf{R}_{2,t}^{-1}\sigma_{n}^{2})}{\partial\mathbf{R}_{2,t}}\!\!=\!\!-\sigma_{n}^{2}\mathbf{R}_{2,t}^{-2}.$ (68)
@@ Appendix A Gradient Derivations
$\frac{\partial\mathrm{KLD}_{\mathrm{r},t}}{\partial\mathbf{W}_{\mathrm{r},t}}=\frac{1}{\ln{2}}(\mathbf{R}_{2,t}^{-1}-\sigma_{n}^{2}\mathbf{R}_{2,t}^{-2})\cdot\frac{\partial\mathbf{R}_{2,t}}{\partial\mathbf{W}_{\mathrm{r},t}},$ (69)
@@
$\!\!\!\!\!\frac{\partial\mathrm{KLD}_{\mathrm{r},t}}{\partial\mathbf{W}_{\!\mathrm{r},t}}\!\!=\!\!\frac{\!d_{\mathrm{r},t}^{\zeta}}{\!L\ln{2}}(\mathbf{H}_{t}^{H}\mathbf{R}_{2,t}^{-1}\mathbf{H}_{t}\!\mathbf{W}_{\!\mathrm{r},t}-\sigma_{n}^{2}\mathbf{H}_{t}^{H}\mathbf{R}_{2,t}^{-2}\mathbf{H}_{t}\!\mathbf{W}_{\!\mathrm{r},t}\!).\!\!$ (70)
@@
$\!\!\!\!\!\![\nabla\!f(\!\mathbf{\dot{W}}\!_{\mathrm{r}}\!)]_{i,l,t}\!\!=\!\!\frac{2d_{\mathrm{r},t}^{\zeta}}{TL\!\ln{2}}(\mathbf{H}_{t}^{H}\mathbf{R}_{2,t}^{-1}\mathbf{H}_{t}\!\mathbf{W}_{\!\mathrm{r},t}\!-\sigma_{n}^{2}\mathbf{R}_{2,t}^{-2}\mathbf{H}_{t}\!\mathbf{W}_{\!\mathrm{r},t})_{i}.\!\!\!$ (71)
@@
$\sigma_{\eta}^{2}=d_{\mathrm{c},k}^{\zeta}\>\sigma_{h}^{2}\mathrm{Tr}(\frac{1}{L}\sum_{l=1}^{L}{\mathbf{W}_{\mathrm{r},l}\mathbf{W}_{\mathrm{r},l}^{H}})+\sigma_{n}^{2}$ (72)
@@
$\frac{\partial\sigma_{\eta}^{2}}{\partial\mathbf{W}_{\mathrm{r},l}}=\frac{2\>d_{\mathrm{c},k}^{\zeta}\sigma_{h}^{2}}{L}\mathbf{W}_{\mathrm{r},l}.$ (73)
@@
$\frac{\partial\mathrm{KLD}_{\mathrm{c},k}}{\partial\sigma_{\eta}^{2}}=-\frac{\lambda\>d_{\mathrm{c},k}^{\zeta}p_{k}(N-K)}{M(M-1)\sigma_{\eta}^{2}\ln 2}.$ (74)
@@
$\frac{\partial\mathrm{KLD}_{\mathrm{c},k}}{\partial\mathbf{W}_{\mathrm{r},l}}=-\frac{2\lambda\>d_{\mathrm{c},k}^{\zeta}p_{k}(N-K)\sigma_{h}^{2}}{M(M-1)L\sigma_{\eta}^{2}\ln 2}\mathbf{W}_{\mathrm{r},l}.$ (75)
@@
$[\nabla\mathrm{KLD}_{\mathrm{c},k}]_{i,l,t}=-\frac{2\lambda\>d_{\mathrm{c},k}^{\zeta}p_{k}(N-K)\sigma_{h}^{2}}{M(M-1)L\sigma_{\eta}^{2}\ln 2}[\mathbf{\dot{W}}_{\mathrm{r}}]_{i,l,t}.$ (76)

### v5

@@ X Conclusion
This paper has proposed and validated a novel KLD-based framework for optimising ISAC systems. The three considered optimisation techniques, namely, K-ROP, K-COP, and K-IOP, each demonstrated significant improvements over both the non-optimised scenario and optimisation benchmarks (D-ROP and B-COP). K-ROP and D-ROP yielded substantial enhancements in target detection and communication performance while K-ROP maintained superior computational efficiency compared to D-ROP, validating the analytical tractability advantages of KLD over traditional Q-function-based BER formulations. K-COP showed notable improvements in the communication subsystem with modest radar gains, particularly at lower $P_{\mathrm{R}}/N_{o}$ levels, and demonstrated more efficient optimisation than B-COP due to KLD’s tractable gradient expressions. The K-IOP approach effectively unified these advantages, consistently outperforming all separate approaches and benchmarks in detection probability, BER, and estimation accuracy. By jointly designing radar and communication waveforms, K-IOP leverages the synergy between sensing and data transmission to achieve higher overall KLD and more balanced performance. The numerical results revealed that exploiting communication signals within the radar subsystem significantly boosts ISAC capabilities, underscoring the importance of holistic design in shared-antenna deployments. These findings provide valuable insights for next-generation ISAC systems, offering a promising framework that adapts system resources for diverse operating regimes while paving the way for more efficient and versatile wireless communications and sensing technologies. Future work will explore the synergistic interplay between KLD and machine learning in ISAC, extend the framework to emerging architectures including cell-free MIMO and full-duplex systems, and develop practical interference exploitation strategies.
This appendix provides detailed derivations for the gradients used in the optimisation problems.
-A Gradients for Radar-Waveform Optimisation
The objective function $f(\mathbf{\dot{W}}_{\mathrm{r}})$ is given in (31), where the $\mathrm{KLD}_{\mathrm{r},t}$ is defined in (27). Therefore, to derive $\nabla f(\mathbf{\dot{W}}_{\mathrm{r}})$, we first compute $\partial\mathrm{KLD}_{\mathrm{r},t}/\partial\mathbf{W}_{\mathrm{r},t}$. Using the matrix identity $\frac{\partial\ln(\det(\mathbf{X}))}{\partial\mathbf{X}}=(\mathbf{X}^{-1})^{T}$, and derivative $\frac{\partial\mathrm{Tr}(\mathbf{X}^{-1}\mathbf{A})}{\partial\mathbf{X}}=-(\mathbf{X}^{-1})^{T}\mathbf{A}^{T}(\mathbf{X}^{-1})^{T}$ we obtain,
$\frac{\partial\ln(\det(\mathbf{R}_{2,t}))}{\partial\mathbf{R}_{2,t}}\!=\!\mathbf{R}_{2,t}^{-1},\quad\frac{\partial\mathrm{Tr}(\mathbf{R}_{2,t}^{-1}\sigma_{n}^{2})}{\partial\mathbf{R}_{2,t}}\!\!=\!\!-\sigma_{n}^{2}\mathbf{R}_{2,t}^{-2}.\vskip-3.61371pt$ (73)
@@
$\frac{\partial\mathrm{KLD}_{\mathrm{r},t}}{\partial\mathbf{W}_{\mathrm{r},t}}=\frac{1}{\ln{2}}(\mathbf{R}_{2,t}^{-1}-\sigma_{n}^{2}\mathbf{R}_{2,t}^{-2})\cdot\frac{\partial\mathbf{R}_{2,t}}{\partial\mathbf{W}_{\mathrm{r},t}},\vskip-3.61371pt$ (74)
@@
$\frac{\partial\mathrm{KLD}_{\mathrm{r},t}}{\partial\mathbf{W}_{\!\mathrm{r},t}}\!\!=\!\!\frac{\!d_{\mathrm{r},t}^{\zeta}}{\!L\ln{2}}(\mathbf{H}_{t}^{H}\mathbf{R}_{2,t}^{-1}\mathbf{H}_{t}\!\mathbf{W}_{\!\mathrm{r},t}-\sigma_{n}^{2}\mathbf{H}_{t}^{H}\mathbf{R}_{2,t}^{-2}\mathbf{H}_{t}\!\mathbf{W}_{\!\mathrm{r},t}\!).\!\!\vskip-3.61371pt$ (75)
@@
$\nabla\!f(\!\mathbf{\dot{W}}\!_{\mathrm{r}}\!)]_{i,l,t}\!\!=\!\!\frac{2d_{\mathrm{r},t}^{\zeta}}{TL\!\ln{2}}(\mathbf{H}_{t}^{H}\mathbf{R}_{2,t}^{-1}\mathbf{H}_{t}\!\mathbf{W}_{\!\mathrm{r},t}\!-\sigma_{n}^{2}\mathbf{R}_{2,t}^{-2}\mathbf{H}_{t}\!\mathbf{W}_{\!\mathrm{r},t})_{i}.\!\!\!\vskip-3.61371pt$ (76)
@@
$\sigma_{\eta}^{2}=d_{\mathrm{c},k}^{\zeta}\>\sigma_{h}^{2}\mathrm{Tr}(\frac{1}{L}\sum_{l=1}^{L}{\mathbf{W}_{\mathrm{r},l}\mathbf{W}_{\mathrm{r},l}^{H}})+\sigma_{n}^{2}\vskip-3.61371pt$ (77)
@@
$\frac{\partial\sigma_{\eta}^{2}}{\partial\mathbf{W}_{\mathrm{r},l}}=\frac{2\>d_{\mathrm{c},k}^{\zeta}\sigma_{h}^{2}}{L}\mathbf{W}_{\mathrm{r},l}.\vskip-3.61371pt$ (78)
@@
$\frac{\partial\mathrm{KLD}_{\mathrm{c},k}}{\partial\sigma_{\eta}^{2}}=-\frac{\lambda\>d_{\mathrm{c},k}^{\zeta}p_{k}(N-K)}{M(M-1)\sigma_{\eta}^{2}\ln 2}.\vskip-3.61371pt$ (79)
@@
$\frac{\partial\mathrm{KLD}_{\mathrm{c},k}}{\partial\mathbf{W}_{\mathrm{r},l}}=-\frac{2\lambda\>d_{\mathrm{c},k}^{\zeta}p_{k}(N-K)\sigma_{h}^{2}}{M(M-1)L\sigma_{\eta}^{2}\ln 2}\mathbf{W}_{\mathrm{r},l}.\vskip-3.61371pt$ (80)
@@
$\nabla\mathrm{KLD}_{\mathrm{c},k}]_{i,l,t}=-\frac{2\lambda\>d_{\mathrm{c},k}^{\zeta}p_{k}(N-K)\sigma_{h}^{2}}{M(M-1)L\sigma_{\eta}^{2}\ln 2}[\mathbf{\dot{W}}_{\mathrm{r}}]_{i,l,t}.\vskip-4.33601pt$ (81)
@@
-B Gradient of $P_{D,t}$ w.r.t. $\mathbf{W}_{\mathrm{r}}$
To enable gradient-based optimisation of $P_{D,t}$, we derive $\nabla_{\mathbf{W}_{\mathrm{r},t}}P_{D,t}$. Using the chain rule with $P_{D,t}=Q(z_{t})$ where $z_{t}\!=\!(\!Q^{-1}(P_{FA})\sigma_{0}\!-\!\Delta\mu\!)/\!\sigma_{1}$ from (43), and $\phi(\!z)\!=\!\frac{1}{\sqrt{\!2\pi}}e^{\!-z^{2}/2}$ being the standard normal PDF, we have,
$\nabla_{\mathbf{W}_{\mathrm{r},t}}P_{D,t}=-\phi(z_{t})\nabla_{\mathbf{W}_{\mathrm{r},t}}z_{t}.\vskip-4.33601pt$ (82)
where $\nabla_{\!\mathbf{W}_{\mathrm{r},t}}z_{t}\!\!=\!\frac{1}{\sigma_{1}}\left(\!Q^{-1}\!(P_{FA})\nabla_{\!\mathbf{W}_{\mathrm{r},t}}\!\sigma_{0}\!-\!\nabla_{\!\mathbf{W}_{\mathrm{r},t}}\!\Delta\mu\!-\!z_{t}\nabla_{\!\mathbf{W}_{\mathrm{r},t}}\!\sigma_{1}\!\right)$. Using the matrix derivative identity $\nabla\!\mathbf{A}^{\!-1}\!=\!-\mathbf{A}^{\!-1}\!(\!\nabla\!\mathbf{A}\!)\mathbf{A}^{\!-1}\!$ obtaining $\mathbf{Q}_{t}$ gradient,
$\nabla_{\mathbf{W}_{\mathrm{r},t}}\mathbf{Q}_{t}\!=\!-\frac{2d_{\mathrm{r},t}^{\zeta}}{L}\mathbf{R}_{1,t}^{-1}\mathbf{H}_{t}\!\mathbf{W}_{\mathrm{r},t}\!\mathbf{W}_{\mathrm{r},t}^{H}\mathbf{H}_{t}^{H}\!\mathbf{R}_{1,t}^{-1}.\vskip-4.33601pt$ (83)
The gradient of $\Delta\mu\!$ using the cyclic property of trace and (83) is,
$\nabla_{\mathbf{W}_{\mathrm{r},t}}\Delta\mu\!=\!\frac{2d_{\mathrm{r},t}^{\zeta}}{L}\mathbf{W}_{\mathrm{r},t}^{H}\mathbf{H}_{t}^{H}[\mathbf{Q}_{t}\!-\!d_{\mathrm{r},t}^{\zeta}\mathbf{R}_{1,t}^{-1}\mathbf{H}_{t}\mathbf{R}_{t}\mathbf{H}_{t}^{H}\mathbf{R}_{1,t}^{-1}]\mathbf{H}_{t}.\vskip-3.61371pt$ (84)
For $\sigma_{0}$ and $\sigma_{1}$, the gradients follow a similar form. The final gradient can be expressed in matrix form as,
$\nabla_{\mathbf{W}_{\mathrm{r},t}}P_{D,t}=-\frac{2d_{\mathrm{r},t}^{\zeta}\phi(z_{t})}{L\sigma_{1}}\mathbf{W}_{\mathrm{r},t}^{H}\mathbf{H}_{t}^{H}\mathbf{M}_{t}\mathbf{H}_{t},\vskip-4.33601pt$ (85)
where $\mathbf{M}_{t}=\mathbf{Q}_{t}-d_{\mathrm{r},t}^{\zeta}\mathbf{R}_{1,t}^{-1}\mathbf{H}_{t}\mathbf{R}_{t}\mathbf{H}_{t}^{H}\mathbf{R}_{1,t}^{-1}+\alpha_{t}\mathbf{Q}_{t}\mathbf{R}_{1,t}^{-1}\mathbf{Q}_{t}\mathbf{R}_{1,t}^{-1}$ with $\alpha_{t}\!=\!(Q^{-1}(\!P_{FA}\!)\sigma_{n}^{2})/(2\sigma_{0}\sigma_{1}\!\sqrt{L})\!-\!z_{t}\sigma_{n}^{2}/\!\sigma_{0}\!\sqrt{L}$.
-C Gradients of BER w.r.t $\mathbf{W}_{\mathrm{c}}$
Using the chain rule, the gradient of $\mathrm{BER}_{k}$ w.r.t. $\mathbf{W}_{\mathrm{c}}$ is:
$\vskip-3.61371pt\nabla_{\mathbf{W}_{\mathrm{c}}}\mathrm{BER}_{k}=\frac{\partial\mathrm{BER}_{k}}{\partial\mathrm{SINR}_{k}}\cdot\nabla_{\mathbf{W}_{\mathrm{c}}}\mathrm{SINR}_{k},\vskip-4.33601pt$ (86)
where the partial derivative is,
$\frac{\partial\mathrm{BER}_{k}}{\partial\mathrm{SINR}_{k}}=-\frac{\lambda_{\mathrm{c}}\sqrt{\beta_{\mathrm{c}}}}{2M\log_{2}M\sqrt{2\pi\mathrm{SINR}_{k}}}e^{-\beta_{c}\mathrm{SINR}_{k}/2},\vskip-4.33601pt$ (87)
and $\nabla_{\mathbf{W}_{\mathrm{c}}}\mathrm{SINR}_{k}$ is derived. Let $\mathcal{D}=d_{\mathrm{c},k}^{-\zeta}\sum_{i\neq k}|\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},i}|^{2}+d_{\mathrm{c},k}^{-\zeta}\|\mathbf{h}_{k}^{T}\mathbf{W}_{\mathrm{r}}\|^{2}+\sigma_{n}^{2}$. Then,
$\nabla_{\mathbf{W}_{\mathrm{c}}}\mathrm{SINR}_{k}$ $=\frac{2d_{\mathrm{c},k}^{-\zeta}}{\mathcal{D}}\mathbf{h}_{k}\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},k}\mathbf{e}_{k}^{T}$
$-\!\frac{2d_{\mathrm{c},k}^{-2\zeta}}{\mathcal{D}^{2}}\|\mathbf{h}_{k}^{T}\!\mathbf{w}_{\mathrm{c},k}\|^{2}\mathbf{h}_{k}\mathbf{h}_{k}^{T}\mathbf{W}_{\!\mathrm{c}}(\mathbf{I}\!-\!\mathbf{e}_{k}\mathbf{e}_{k}^{T}),\vskip-4.33601pt$ (88)
where $\mathbf{e}_{k}$ $k$-th standard basis vector. Substituting (87) and (88) into (86), we get the gradient $\nabla_{\mathbf{W}_{\mathrm{c}}}\mathrm{BER}_{k}$.